answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

In a study using 9 samples and in which the population variance is unknown the distribution that should be used to calculate confidence intervals is?

In a study using 9 samples, and in which the population variance is unknown, the distribution that should be used to calculate confidence intervals is

Is 13 a Fibonacci number?

the first seven Fibonacci numbers are 1,1,2,3,5,8,13. 13 is a Fibonacci number.

A bag contains 9 marbles 2 green 3 blue and 4 red what is the pobability of picking a green then a non-green?

Short answer: 14/81

There is a 2/9 chance of picking a green marble the first time.

Since this leaves 8 marbles (only one of which is green; you took the other one), that leaves an 7/8 chance of picking any other colour the second time.

To find the overall probability, you simply multiply the two values (2/9 * 7/8 = 14/81). This is simple to understand. You can only get a green marble every 2 out of 9 tries. For each of these 2 times, you have a 7/8 chance of getting a different colour the second time. So, you have a 7/8 chance every 2/9 times. Simply multiply them to get the total number of possibilities.

You roll a number cube 150 times about how many times would you expect to roll 3?

Possible outcomes of one roll = 6

Successful outcomes = 1

Probability of success on each roll = 1/6

Expectation in 150 rolls = (1/6) x (150) = 25 times

What does less probable mean?

n event A is said to be less probable than an event B if the probability of A occurring is smaller than that of B occurring.

Why is the sample mean an unbiased estimator of the population mean?

The sample mean is an unbiased estimator of the population mean because the average of all the possible sample means of size n is equal to the population mean.

If you have a bag containing 24 marbles If the probability of drawing a red marble is twice as likely not drawing a red marble how many red marbles are in the bag?

Suppose probability of drawing a red marble is p.

Then p = 2*(1 - p)

that is p = 2 - 2p

or p = 2/3

So 2/3 of the 24 marbles are red

24*(2/3) = 16 red marbles.

Probability lesson represent outcomes 4th grade math?

That's not a question; indeed, it isn't even a sentence. If you're looking for a lesson to teach to 4th-graders, you might try an educators' site instead.

In which discrete distributions does the probability of a success vary from one trial to the next?

In situations like this it's more common to model the systems as what are called stochastic processes, in this case probably Markov chains.

What effect does sample size have on the estimate of the mean?

the larger you r sample size the better your estimate. imagine take the height of person to estimate the average high of an adult male. would one person's height be a good estimate, or would taking the average height of 100, or 5000 adult males will produce a better estimate?

Theoretical probability of rolling a factor of 10 on a number cube?

The factors of 10 are the numbers that divide 10 evenly: 1, 2, 5 and 10. To answer your question, you have to figure out what the probability of rolling one of these numbers is on a number cube.

When do you use uniform distribution?

When, over a given range, the probability that a variable in question lies within a particulat interval is equal to the size of that interval as a proportion of the range.

In statistics which is better large sample size or small sample size?

Generally, the larger the sample the more reliable the results.

Example:

If you flipped a coin twice and got heads both times you could say the coined is biased towards heads.

However, if you repeat the experiment 100 times your results will be a lot more reliable.

What do you win if the odds are 5 to 2?

When the odds are 5 to 2, if you bet $2 on something and win then you will receive $7 The $2 you bet and the $5 you won.

Juan rolled two fair number cubes that have faces labeled 1 to 6 each he rolled the two cubes 210 times according to theoretical probability how many of the rolls should juan expect the sum to be 7?

possible ways to get 7:

1+6=7

6+1=7

2+5=7

5+2=7

3+4=7

4+3=7

6 possible ways to get 7. There are 36 possible ways the two dice could be rolled, so you would get 6/36 or in this case, 35/210

Answer:35

What is the probability that a couple would have a family of four children two boys and two girls and in that order?

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a boy is approx 0.52 in all cases, the overall probability is 0.0624.

Trending Questions
possible outcomes if two dices X and Y are rolled simultaneously and make a list of the different outcomes of which the total is 6? What is the duce in a pack of playing cards? Is a bell shaped curve an example of a skewed distribution? What is the probability of drawing two 10s in card deck? Why the sample should be normally distributed? What is the probability of drawing a red king from a deck of 52? Is ace before 2? How many red face cards in a deck of playing cards? What is the probability it will occur equals 0? A bag contains 6 red marbles 9 blue marbles and 5 green marbles You withdraw one marble replace it and withdraw another marble What is the probability that you do not draw two green marbles 25 denomo? Is this independent or dependent p (drawing a face card from a deck of cards after a jack was just drawnand replaced and the deck shuffled again? Which type of writing is best illustrated in a Venn diagram? How can risk be assessed during an IS justification process? How many ways can the word Tennessee be arranged? What is selecting customers? What instructions is a patient given after undergoing a D and C? What is the formula for standard error? When a single card is drawn drawn from an ordinary 52 card deck find the probability of getting either a king or a queen? What is the probability of the following events P(b and an even number)? What is the standard error of the sampling distribution equal to when you do not know the population standard deviation?